OpenStudy (anonymous):

what is the value of 238 - 3x^2 when x = 4?

OpenStudy (popitree):

238-3(4^2)

OpenStudy (anonymous):

-12

OpenStudy (anonymous):

A: 226 B: 222 C: 190 D: 110

OpenStudy (anonymous):

sorry mistype C

OpenStudy (imtiaz7):

190 is right ans

OpenStudy (popitree):

190

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

i got another one if your up fo it

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

what is the value of 8x^2 - 5x when x = 3

OpenStudy (anonymous):

A: 96 B: 57 C: 33 D:19

OpenStudy (anonymous):

a

OpenStudy (anonymous):

thanks bro, one more? Last one

OpenStudy (anonymous):

what is 56 / 7 + 3y^3 when y = 4

OpenStudy (anonymous):

A: 200 B: 156 C: 132 D: 44

OpenStudy (popitree):

what is the value of 8x^2 - 5x when x = 3 == 57, not 96

OpenStudy (anonymous):

c

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

jman C for 8x^2?

OpenStudy (imtiaz7):

A 200 is right one :)